OpenStudy (anonymous):

The light turns green, and a bicyclist starts forward with an acceleration of 1.2m/s^2 . How far must she travel to reach a speed of 8.0m/s ?

mathslover (mathslover):

given a = 1.2 m/s^2 and also u = 0 m/s and v = 8.0 m/s

mathslover (mathslover):

\[\large{v^2-u^2=2as}\]

OpenStudy (anonymous):

so 26.7?

mathslover (mathslover):

\[\large{64 = 2(1.2) s}\] \[\large{64=2.4s}\] \[\large{\frac{64}{2.4}=s}\] \[\large{\frac{640}{24}=s}\] \[\large{26.7 m = s}\]

mathslover (mathslover):

good work @mandonut

OpenStudy (anonymous):

i used the other formula

OpenStudy (anonymous):

with at^2

mathslover (mathslover):

you can use any formula but acc. to me this is easier

OpenStudy (anonymous):

okay thanks

OpenStudy (lgbasallote):

hmm weird..usually math slover goes on saying how these are physics questions...wonder why the change of heart

mathslover (mathslover):

for that u must apply v = u + at first and then use the value of t

mathslover (mathslover):

@lgbasallote that is because it includes the work of math and yes I was tired saying that all .. at last users continue that they want to do... so no response here

OpenStudy (lgbasallote):

ahh welcome to the dark side

mathslover (mathslover):

:D but this will not be in dark side..when I will recover myself with sickness